OpenStudy (anonymous):

Evaluate the integral, where E is the solid in the first octant that lies beneath the paraboloid z = 1 - x2 - y2. Use cylindrical coordinates. triple integral (4(x^3+xy^3)) dV

OpenStudy (anonymous):

z is going to range from: \[0\le z \le 1-r^2\]. Then let z=0 Solving you get: \[x^2+y^2=1\]. Which implies: \[0 \le r \le 1\]. And its in the first octant so: \[0 \le \theta \le \frac{\pi}{2}\]. Just don't forget when you do dV to make it: \[r dz dr d \theta\]

OpenStudy (anonymous):

Then you need to convert your integrand: \[4(x^3+xy^3)=4(r^3\cos^3(\theta)+r^4\cos(\theta)\sin^3(\theta))\]

OpenStudy (anonymous):

So you should have: \[4\int\limits_{0}^{\frac{\pi}{2}} \int\limits_{0}^{1} \int\limits_{0}^{1-r^2}(r^3\cos^3(\theta)+r^4\cos(\theta)\sin^3(\theta))rdzdrd \theta\]. Not pretty but should be doable :P